# The Unbounded Denominators Conjecture

### v1

@@ 5. The uniformization of $\mathbf{C}\smallsetminus\mu_{N}$
In this section we develop all the particular analytic properties that we need of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$. André has pointed out to us that our two main results here, Theorem 5.3.8 and Lemma 5.5.7, appear in work of Kraus and Roth [KR16, Remark 5.1 and Theorems 1.2 and 1.10]. Nevertheless, as our proofs are simplified to cover our current needs, and since the results of Kraus and Roth rely on some previous work of themselves and others, we keep our self-contained exposition as a convenience to the reader, and refer to [ASVV10, KRS11, KR16] and the references there for various further results and a more thorough study of the uniformization of $\mathbf{C}\smallsetminus\mu_{N}$.
@@
This allows one to pass freely between uniformizations by $D(0,1)$ and $\mathbf{H}$. In this section, we choose notation so that the corresponding passage from $D(0,1)$ to $\mathbf{H}$ is marked by the addition of a tilde. Thus, for example, $\widetilde{F}_{N}$ constructed below denotes a map on $\mathbf{H}$ and $F_{N}$ (Definition 5.1.1) is simply the pull-back of $\widetilde{F}_{N}$ to $D(0,1)$ via the map above. Similarly, $\Gamma_{N}$ will denote a lattice in $U(1,1)$ whereas $\widetilde{\Gamma}_{N}$ denotes the corresponding lattice in $\mathrm{PSL}_{2}(\mathbf{R})$.
@@
$\widetilde{F}_{N}:\mathbf{H}\rightarrow\mathbf{H}/\widetilde{\Gamma}_{N}=\mathbf{C}\smallsetminus\mu_{N}.$
The map $\widetilde{F}_{N}$ is unique up to the action of $\mathrm{PSL}_{2}(\mathbf{R})$ on the source — we pin down some precise choices in §5.1, but the analysis of this subsection will not depend on any such choices.
@@
We have a companion uniformization map $1/\widetilde{F}_{N}:\mathbf{H}\rightarrow\mathbf{P}^{1}\smallsetminus\{0,\mu_{N}\}$. We now derive the Schwarzian derivative of the inverse map (of $1/\widetilde{F}_{N}$ and then of $\widetilde{F}_{N}$) following [Hem88]. (The reason for first considering the reciprocal of $\widetilde{F}_{N}$ is that the standard form considered in [Hem88] is for maps to $\mathbf{P}^{1}\smallsetminus S$ where $S$ is a finite set of points which does not contain $\infty$.) Let $m_{k}$ for $k=1\ldots N$ denote the accessory parameters at $z=p_{k}=\zeta^{k}=e^{2\pi ik/N}$, and let $m_{0}$ denote the accessory parameter at $z=p_{0}=0$. Exactly as in [Hem88, Example 1], the accessory parameters $m_{k}$ for $k\neq 0$ satisfy the symmetry $m_{k}=c\cdot\zeta^{-k}$ for some constant $c$. The accessory parameters are subject to three constraints, given by ([Hem88, Theorem 3.1])
@@
From the first constraint, we deduce from the fact that $\sum_{k=1}^{N}m_{k}=0$ that $m_{0}=0$. But now from the second constraint in equation (5.0.2), we deduce that
$\sum_{k=0}^{N}(2m_{k}\zeta^{k}+1)=1+\sum_{k=1}^{N}(2c+1)=0$
and hence $c=-\frac{1}{2}-\frac{1}{2N}$. Writing $Z=1/F_{N}(\tau)$, we deduce from [Hem88, Theorem 3.1] that the Schwarzian $\{\tau,1/F_{N}\}$ is given by
$\{\tau,1/F_{N}\}=\{\tau(Z),Z\}:=\left(\Big(\frac{\tau^{\prime\prime}}{\tau^{\prime}}\Big)^{\prime}-\frac{1}{2}\left(\frac{\tau^{\prime\prime}}{\tau^{\prime}}\right)^{2}\right)(Z)=\frac{1}{2}\sum_{k=0}^{N}\frac{1}{(Z-p_{k})^{2}}+\sum_{k=0}^{N}\frac{m_{k}}{Z-p_{k}}$
$\frac{1}{2Z^{2}}+\frac{1}{2}\sum_{k=1}^{N}\frac{1}{(Z-\zeta^{k})^{2}}-\frac{(1+N)}{2N}\sum_{k=1}^{N}\frac{\zeta^{-k}}{Z-\zeta^{k}}=\frac{(1+(N^{2}-1)Z^{N})}{2Z^{2}(Z^{N}-1)^{2}}.$
From the chain rule, we deduce that with $z=F_{N}=1/Z$ the equality:
$\{\tau,F_{N}\}=\frac{1}{z^{4}}\frac{(1+(N^{2}-1)(1/z)^{N})}{2(1/z)^{2}((1/z)^{N}-1)^{2}}=\frac{(N^{2}-1)z^{N-2}+z^{2N-2}}{2(z^{N}-1)^{2}}.$
By [Hem88, Lemma 3.3], if $\eta_{1}$ and $\eta_{2}$ are solutions to the equation $y^{\prime\prime}+\frac{1}{2}\{\tau,F_{N}\}y=0$, or equivalently
(5.0.3) $4(z^{N}-1)^{2}y^{\prime\prime}+((N^{2}-1)z^{N-2}+z^{2N-2})y=0$
then the (locally analytic) inverse map $\psi_{N}$ of $\widetilde{F}$ is the ratio $\eta_{1}/\eta_{2}$ up to a Möbius transformation (equivalently, up to the correct choice of linearly independent solutions). We find that equation (5.0.3) admits solutions
(5.0.4) $\eta_{1}=1-\frac{(N+1)z^{N}}{4N}+\cdots\in\mathbf{Q}\llbracket z^{N}\rrbracket,$
(5.0.5) $\eta_{2}=z-\frac{(N-1)z^{N+1}}{4N}+\cdots\in z\cdot\mathbf{Q}\llbracket z^{N}\rrbracket,$
the inclusions following in an elementary way from the Frobenius method applied to (5.0.3).
5.1. Normalizations of $\widetilde{F}_{N}$
Using the action of $\mathrm{PSL}_{2}(\mathbf{R})$ we may assume that $\widetilde{F}_{N}(i\infty)=1$. The stabilizer of $\infty$ consists of Möbius transformations $z\mapsto az+b$, so by specifying $\widetilde{F}_{N}(0)=0$ we determine $\widetilde{F}_{N}$ uniquely. But note that if $\zeta_{N}=\exp(2\pi i/N)$, then $\zeta_{N}\widetilde{F}_{N}$ is another covering map which must thereby differ from $\widetilde{F}_{N}$ by a Möbius transformation $\zeta_{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(r_{N}\cdot\tau)$ for of $i\in\mathbf{H}$. We deduce that $\widetilde{F}_{N}(r^{N}_{N}\cdot\tau)=\zeta^{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\tau)$, and thus $r^{N}_{N}\in\mathrm{SO}_{2}(\mathbf{R})$ must also lie in $\widetilde{\Gamma}_{N}$. But $\widetilde{\Gamma}_{N}$ is a free group, and hence $r^{N}_{N}$ is trivial, and $r_{N}$ is a hyperbolic rotation around $i$ of order $N$.
@@ 5.1. Normalizations of $\widetilde{F}_{N}$
Let us define $F_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\mu_{N}$ by the formula
$F_{N}(x)=\widetilde{F}_{N}\left(i\cdot\frac{1+x}{1-x}\right).$
Note that this is just the map $\widetilde{F}_{N}$ composed with the standard conformal isomorphism $D(0,1)\rightarrow\mathbf{H}$ sending 0 to $i$, and hence
$F_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\mu_{N}$
is a universal covering map. The action of $\mathrm{SO}_{2}(\mathbf{R})$ on $D(0,1)$ under the pullback map is just given by rotation, and hence $r_{N}$ acts on $D(0,1)$ by a rotation of order $N$. We deduce that $F_{N}(\zeta^{m}q)=\zeta F(q)$ for some $(m,N)=1$. Since $F_{N}$ is a covering map and $F_{N}(0)=0$, we must also have $F^{\prime}_{N}(0)\neq 0$, and thus $F_{N}(\zeta q)=\zeta F_{N}(q)$ for any $N$th root of unity $\zeta$.
Definition 5.1.2.
Let $\psi_{N}$ be the the inverse of $F_{N}$.
The map $\psi_{N}$ is well-defined up to the action of $\Gamma_{N}\subset\mathrm{PSU}(1,1)$ acting on $D(0,1)$. in a neighbourhood of $q=0$ so $\psi_{N}(0)=0$. If $\widetilde{\psi}_{N}$ is the local inverse of $F_{N}$ around 0 so that $\widetilde{\psi}_{N}(0)=i$, then clearly $\psi_{N}(z)$ is a Möbius translate of $\psi_{N}$ and hence also of $\eta_{2}/\eta_{1}$. But now from the equality $F_{N}(\zeta q)=\zeta F_{N}(q)$ we deduce (from the form of $\eta_{1}$ and $\eta_{2}$ in equations (5.0.4) and (5.0.5)) that
(5.1.3) $\psi_{N}(z)=\gamma^{-1}_{N}\cdot\frac{\eta_{2}}{\eta_{1}}=\gamma^{-1}_{N}\left(z+\frac{z^{N+1}}{2N}+\cdots\right)$
for some constant $\gamma^{-1}_{N}$. Moreover, we certainly also have
$F_{N}(q)=\gamma_{N}q+O(q^{2}),$
and hence the conformal radius of $F_{N}$ is given by $|\gamma_{N}|$. We shall compute $\gamma_{N}$ explicitly in Theorem 5.3.8 below.
Definition 5.1.4.
Let $G_{N}$ denote the map $D(0,1)\rightarrow\mathbf{C}\smallsetminus\{1\}$ such that $G_{N}(q^{N})=(F_{N}(q))^{N}$, or equivalently $G_{N}(q)=(F_{N}(q^{1/N}))^{N}$.
The fact that $G_{N}$ is well-defined is a formal consequence of the relation $F_{N}(\zeta q)=\zeta F_{N}(q)$. From the fact that $F_{N}$ is the universal covering map, we also immediately deduce the following geometric description of $G_{N}$.
@@
The map $G_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\{1\}$ is a covering map away from the point $0\in\mathbf{C}\smallsetminus\{1\}$. The map $G_{N}$ is locally an isomorphism in a neighbourhood of 0 in $D(0,1)$, but it is totally ramified of degree $N$ at all other preimages of 0. Moreover, it is universal with respect to any such map.
The inverse map of $G_{N}$ is closely related to the inverse map of $F_{N}$, and turns out to have a nicer form:
Lemma 5.1.6.
Let $\varphi_{N}$ denote the inverse map of $G_{N}$, normalized so that $\varphi_{N}(0)=0$. The function $\varphi_{N}$ has the form $\delta^{-1}_{N}(\phi_{2}/\phi_{1})^{N}$, where $\phi_{1}$ and $\phi_{2}$ are the solutions to the differential equation:
@@
such that
(5.1.8) $\phi_{1}=1-\frac{(N+1)z}{4N}+\cdots,\qquad\phi_{2}=z^{1/N}\left(1-\frac{(N-1)z}{4N}+\cdots.\right).$
The conformal radius of the map $G_{N}$ is $|\gamma^{N}_{N}|$.

### v2

@@ 5. The uniformization of $\mathbf{C}\smallsetminus\mu_{N}$
In this section we develop all the particular analytic properties that we need of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ for $N\geq 2$. André has pointed out to us that our two main results here, Theorem 5.1.4 and Lemma 5.2.17, appear in work of Kraus and Roth [KR16, Remark 5.1 and Theorems 1.2 and 1.10]. Nevertheless, as our proofs are simplified to cover our current needs, and since the results of Kraus and Roth rely on some previous work of themselves and others, we keep our self-contained exposition as a convenience to the reader, and refer to [ASVV10, KRS11, KR16] and the references there for various further results and a more thorough study of the uniformization of $\mathbf{C}\smallsetminus\mu_{N}$. The reader will also benefit from the material in § III.1 in Goluzin’s book [Gol69], which recovers $F_{N}$ via an explicit computation of the Riemann map of a $\mathbf{Z}/N\mathbf{Z}$-rotationally symmetric circular $N$-gon, taking the case of zero angles and doing Schwarz reflections in the sides of the circular polygon.
@@
This allows one to pass freely between uniformizations by $D(0,1)$ and $\mathbf{H}$. In this section, we choose notation so that the corresponding passage from $D(0,1)$ to $\mathbf{H}$ is marked by the addition of a tilde. Thus, for example, $\widetilde{F}_{N}$ constructed below denotes a map on $\mathbf{H}$ and $F_{N}$ (Definition 5.1.1) is simply the pull-back of $\widetilde{F}_{N}$ to $D(0,1)$ via the map above. Similarly, $\Gamma_{N}$ will denote a lattice in $\mathrm{PSU}(1,1)$ whereas $\widetilde{\Gamma}_{N}$ denotes the corresponding lattice in $\mathrm{PSL}_{2}(\mathbf{R})$.
Unless we expressly state otherwise, we reserve $z,\tau,x$ to denote the coordinates on $\mathbf{C}\smallsetminus\mu_{N}$ (or $\mathbf{C}\smallsetminus\{1\}$), $\mathbf{H}$, and $D(0,1)$ respectively.
5.1. Schwarzians and the conformal radius
@@
$\widetilde{F}_{N}:\mathbf{H}\rightarrow\mathbf{H}/\widetilde{\Gamma}_{N}=\mathbf{C}\smallsetminus\mu_{N},$
where $\widetilde{\Gamma}_{N}\subset\mathrm{PSL}_{2}(\mathbf{R})$ denotes the Fuchsian group of $\mathbf{C}\smallsetminus\mu_{N}$. The map $\widetilde{F}_{N}$ is unique up to the action of $\mathrm{PSL}_{2}(\mathbf{R})$ on the source.
We now pin down the definition of $\widetilde{F}_{N}$ with the following precise choices. Using the action of $\mathrm{PSL}_{2}(\mathbf{R})$ we may assume that $\widetilde{F}_{N}(i)=0$. The stabilizer of $i\in\mathbf{H}$ consists of Möbius transformations given by elements in $\mathrm{SO}_{2}(\mathbf{R})$, so by specifying $\widetilde{F}_{N}(i\infty)=1$ (here by abuse of notation, we also use $\widetilde{F}_{N}$ to denote the natural extension of $\widetilde{F}_{N}$ on $\mathbf{H}$ to cusps), we determine $\widetilde{F}_{N}$ uniquely.
Definition 5.1.1.
Define $F_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\mu_{N}$ by the formula
$F_{N}(x)=\widetilde{F}_{N}\left(i\cdot\frac{1+x}{1-x}\right).$
Note that $F_{N}$ is just the map $\widetilde{F}_{N}$ composed with the standard conformal isomorphism $D(0,1)\rightarrow\mathbf{H}$ sending 0 to $i$, and hence
$F_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\mu_{N}$
is the universal covering map with $F_{N}(0)=0$ and $F_{N}(1)=1$.
Note that the statements of the main results of this section, Theorem 5.1.4 and Lemma 5.2.17, only depend on the normalization $F_{N}(0)=0$ and do not depend on the choice $F_{N}(1)=1$.
The following lemma gives the basic symmetric property of $\widetilde{F}_{N}$ and $F_{N}$.
Lemma 5.1.2.
Let $\zeta_{N}=\exp(2\pi i/N)$ and $\zeta$ be any $N$th root of unity. Then $\zeta_{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\widetilde{r}_{N}\cdot\tau)$ and $F_{N}(\zeta x)=\zeta F_{N}(x)$, where
(5.1.3) $\widetilde{r}_{N}=\left(\begin{matrix}\cos(\pi/N)&-\sin(\pi/N)\\ \sin(\pi/N)&\cos(\pi/N)\end{matrix}\right)\in\mathrm{SO}_{2}(\mathbf{R}).$
Proof.
Note that $\zeta_{N}\widetilde{F}_{N}$ is another covering map such that $\zeta_{N}\widetilde{F}_{N}(i)=0$. Therefore $\zeta_{N}\widetilde{F}_{N}$ must differ from $\widetilde{F}_{N}$ by a Möbius transformation in the stabilizer of $i$; that is $\zeta_{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\widetilde{r}_{N}\cdot\tau)$ for some $\widetilde{r}_{N}\in\mathrm{SO}_{2}(\mathbf{R})$. We deduce that $\widetilde{F}_{N}(\widetilde{r}^{N}_{N}\cdot\tau)=\zeta_{N}^{N}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\tau)$, and thus $\widetilde{r}^{N}_{N}\in\mathrm{SO}_{2}(\mathbf{R})$ must also lie in $\widetilde{\Gamma}_{N}$. But $\widetilde{\Gamma}_{N}$ is a free group (due to the fact that $\widetilde{F}_{N}$ is a covering map with no ramification points), and hence $\widetilde{r}^{N}_{N}$ is trivial in $\mathrm{SO}_{2}(\mathbf{R})/\{\pm I\}$, and $\widetilde{r}_{N}$ is a hyperbolic rotation around $i$ of order $N$.
The action of $\mathrm{SO}_{2}(\mathbf{R})$ on $D(0,1)$ under the pullback map is just given by rotation, and hence $\widetilde{r}_{N}$ acts on $D(0,1)$ by a rotation of order $N$. We deduce that $F_{N}(\zeta^{m}x)=\zeta F_{N}(x)$ for some $(m,N)=1$. By taking the derivatives with respect to $q$ of both sides at $q=0$, we have $\zeta^{m}F_{N}^{\prime}(0)=\zeta F_{N}^{\prime}(0)$. Since $F_{N}$ is a covering map, we must also have $F^{\prime}_{N}(0)\neq 0$, and thus $m=1$ and $F_{N}(\zeta x)=\zeta F_{N}(x)$. We thus also deduce (5.1.3) as this action corresponds to the rotation by degree $2\pi/N$ on $D(0,1)$. ∎
Our first main goal of this section is an explicit computation of the uniformization radius of $\mathbf{C}\smallsetminus\mu_{N}$. This formula has been previously proved by Kraus and Roth in [KR16, Remark 5.1].
Theorem 5.1.4.
The conformal size $|F_{N}^{\prime}(0)|$ (Riemann uniformization radius of $\mathbf{C}\smallsetminus\mu_{N}$) is equal to
(5.1.5) $|F_{N}^{\prime}(0)|=\gamma_{N}:=16^{1/N}\frac{\displaystyle{\Gamma\left(1+\frac{1}{2N}\right)^{2}\Gamma\left(1-\frac{1}{N}\right)}}{\displaystyle{\Gamma\left(1-\frac{1}{2N}\right)^{2}\Gamma\left(1+\frac{1}{N}\right)}}.$
We have an expansion for $\gamma_{N}$ as follows:
(5.1.6) $\gamma_{N}=16^{1/N}\left(1+\frac{\zeta(3)}{2N^{3}}+\frac{3\zeta(5)}{8N^{5}}+O(N^{-6})\right),$
where the remaining term $O(N^{-6})$ is a positive real number.
To prove this formula, we follow Hempel [Hem88] to get a second order linear ODE whose ratio of two linearly independent solutions gives the (local analytic) inverse of $\widetilde{F}_{N}$ (Lemma 5.1.8). The uniformization maps of Riemann surfaces—and their inverses—do not typically admit explicit solutions in terms of standard functions, but our particular case of interest turns out to be an exception due to the extra symmetries of $\mathbf{C}\smallsetminus\mu_{N}$. We use Lemma 5.1.2 to define a function $G_{N}$ closely related to $F_{N}$ (see Definition 5.1.13) and explicitly find two solutions of the associated linear ODE in terms of hypergeometric functions (Lemma 5.1.14). These solutions allow us to compute the explicit conformal radius for $G_{N}$ and deduce the formula (5.1.5) for $F_{N}$.
Our computation here is very similar to the treatment by Goluzin in [Gol69, §III.1], who also gives the explicit formula for the inverse of $G_{N}$ by hypergeometric functions. See the $q=0$ case of equation (17) and the last paragraph on page 86 of loc. cit. Goluzin more generally computes the Riemann map for the $\mathbf{Z}/N\mathbf{Z}$-rotationally symmetric circular $N$-gon with angles $\pi q$, and explains [Gol69, § II.6] how the $q=0$ case (formula (21) on page 86 of loc. cit.) by Schwarz reflections entails a description of $G_{N}$ and $F_{N}$.
Definition 5.1.7.
Let $\psi_{N}$ be the local analytic inverse of $F_{N}$ such that $\psi_{N}(0)=0$.
This inverse exists and is unique in a small neighborhood of $z=0$. As all we need is to compute $F^{\prime}_{N}(0)=\psi_{N}^{\prime}(0)^{-1}$, having $\psi_{N}$ well-defined in a small neighborhood of $z=0$ is enough for our purpose.
Lemma 5.1.8.
The local analytic inverse map $\psi_{N}$ of $F_{N}$ has the form $\psi_{N}=\eta_{1}/\eta_{2}$, where $\eta_{1}$ and $\eta_{2}$ satisfy the second order linear differential equation
(5.1.9) $4(z^{N}-1)^{2}y^{\prime\prime}+((N^{2}-1)z^{N-2}+z^{2N-2})y=0.$
Remark 5.1.10.
The equation (5.1.9) is more transparent in terms of the Schwarzian derivative $\{\tau,\widetilde{F}_{N}\}$ defined and computed in the proof below:
$y^{\prime\prime}+\frac{1}{2}\{\tau,\widetilde{F}_{N}\}y=0.$
Proof.
@@
First, since $F_{N}$ is the composition of $\widetilde{F}_{N}$ and a Möbius transformation, we only need to prove the similar assertion for $\widetilde{F}_{N}$. We have a companion uniformization map $1/\widetilde{F}_{N}:\mathbf{H}\rightarrow\mathbf{P}^{1}\smallsetminus\{0,\mu_{N}\}$. (The reason for first considering the reciprocal of $\widetilde{F}_{N}$ is that the standard form considered in [Hem88] is for maps from $\mathbf{H}$ to $\mathbf{P}^{1}\smallsetminus S$ where $S$ is a finite set of points which does not contain $\infty$.)
[…]
We now derive $\{\tau,1/\widetilde{F}_{N}\}$ and then $\{\tau,\widetilde{F}_{N}\}$ following [Hem88, §3, §6].
Let $p_{k}=\zeta_{N}^{k}=e^{2\pi ik/N}$ for $k=1,\ldots,N$ and let $p_{0}=0$. We deduce from [Hem88, Theorem 3.1] that the Schwarzian $\{\tau,1/\widetilde{F}_{N}\}$ is given by
(5.1.11) $\{\tau,1/\widetilde{F}_{N}\}=\frac{1}{2}\sum_{k=0}^{N}\frac{1}{(Z-p_{k})^{2}}+\sum_{k=0}^{N}\frac{m_{k}}{Z-p_{k}},$
where the $m_{k}$ for $k=0,\ldots,N$ denote the so-called accessory parameters at $z=p_{k}$. The accessory parameters are notoriously hard to compute in general, but in our particular example we may find them using the $\mathbf{Z}/N\mathbf{Z}$ symmetry. Expressing the fact that (5.1.11) vanishes to order four at $Z=\infty$, the accessory parameters are subject the following three constraints [Hem88, Theorem 3.1] obtained by equating the $1/Z,1/Z^{2}$ and $1/Z^{3}$ coefficients to zero:
@@
Since $\mathbf{C}\smallsetminus\mu_{N}$ is invariant under the action of $\mu_{N}$, we deduce exactly as in [Hem88, §6, Example 1] that the accessory parameters $m_{k}$ for $k\neq 0$ satisfy the symmetry $m_{k}=c\cdot\zeta_{N}^{-k}$ for some constant $c$. The constraint $\sum_{k=1}^{N}m_{k}=0$ in (5.1.12) then gives $m_{0}=0$. Then the second constraint in (5.1.12) gives
$\sum_{k=0}^{N}(2m_{k}\zeta_{N}^{k}+1)=1+\sum_{k=1}^{N}(2c+1)=0,$
and hence $c=-\frac{1}{2}-\frac{1}{2N}$. This determines all the $m_{k}$, and turns (5.1.11) (still with $Z=1/\widetilde{F}_{N}$) into
$\{\tau,1/\widetilde{F}_{N}\}=\frac{1}{2Z^{2}}+\frac{1}{2}\sum_{k=1}^{N}\frac{1}{(Z-\zeta_{N}^{k})^{2}}-\frac{(1+N)}{2N}\sum_{k=1}^{N}\frac{\zeta_{N}^{-k}}{Z-\zeta_{N}^{k}}=\frac{(1+(N^{2}-1)Z^{N})}{2Z^{2}(Z^{N}-1)^{2}}.$
From the chain rule, we deduce that with $z=\widetilde{F}_{N}=1/Z$ the equality:
$\{\tau,\widetilde{F}_{N}\}=\frac{1}{z^{4}}\frac{(1+(N^{2}-1)(1/z)^{N})}{2(1/z)^{2}((1/z)^{N}-1)^{2}}=\frac{(N^{2}-1)z^{N-2}+z^{2N-2}}{2(z^{N}-1)^{2}},$
and from this we find that the equation $y^{\prime\prime}+\frac{1}{2}\{\tau,\widetilde{F}_{N}\}y=0$ is given by (5.1.9). We then conclude the proof by [Hem88, Lemma 3.3]. ∎
@@
Let $G_{N}$ denote the map $D(0,1)\rightarrow\mathbf{C}\smallsetminus\{1\}$ such that $G_{N}(x^{N})=(F_{N}(x))^{N}$, or equivalently $G_{N}(x)=(F_{N}(x^{1/N}))^{N}$.
The fact that $G_{N}$ is well-defined is a formal consequence of the relation $F_{N}(\zeta x)=\zeta F_{N}(x)$ in Lemma 5.1.2.
The inverse map of $G_{N}$ is closely related to the inverse map of $F_{N}$, and turns out to have a nicer form. We will give some geometric description of $G_{N}$ in §5.2 in terms of triangle groups, which suggests an explicit description of the inverse of $G_{N}$ in terms of hypergeometric functions.
@@
Let $\varphi_{N}$ denote the local inverse map of $G_{N}$ around $z=0$, normalized so that $\varphi_{N}(0)=0$. The function $\varphi_{N}$ has the form $\delta^{-1}_{N}(\phi_{1}/\phi_{2})^{N}$, where $\phi_{1}$ and $\phi_{2}$ are the solutions to the differential equation:
@@
given explicitly by
[…]
and $\delta_{N}=|G_{N}^{\prime}(0)|$ denotes conformal radius of the map $G_{N}$.
Moreover, let $s_{N}(z)$ denote the function
[…]
Then $\varphi_{N}(z)=\delta_{N}^{-1}s_{N}(z)^{N}$, $\psi_{N}(z)=|F^{\prime}_{N}(0)|^{-1}s_{N}(z^{N})$, and $\delta_{N}=|F^{\prime}_{N}(0)|^{N}$.